# Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning

### v2

@@ A-B Proof of Transfer Sensitivity
Let $\mathcal{G}=\nabla_{\theta}\mathcal{L}(\theta\mid M)$ and recall
@@
We model perturbations around a single update step as
@@
Assume $L(\theta\mid M)$ is differentiable and locally $\beta$-smooth around $\theta$, i.e., $\|\nabla\mathcal{L}(\theta+\delta\mid M)-\nabla\mathcal{L}(\theta\mid M)\|_{2}\leq\beta\|\delta\|_{2}$. Then a first-order Taylor expansion yields
@@
and hence
$S(\theta,\delta\mid M)=-\mathcal{G}^{\top}\delta-r(\delta).$ (21)
Therefore,
$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)=-\mathcal{G}^{\top}(\delta-\delta^{\prime})-(r(\delta)-r(\delta^{\prime})).$ (22)
When perturbations are sufficiently small (so the remainder term is dominated), the leading term is
$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)\approx\mathcal{G}^{\top}(\delta^{\prime}-\delta)=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi).$ (23)
@@
Let $X:=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi)$. Since $\xi^{\prime}-\xi\sim\mathcal{N}(0,2\sigma^{2}I)$, we have $X\sim\mathcal{N}\!\big(0,\,2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}\big)$ and thus
$\mathbb{E}|X|\leq\sqrt{\mathbb{E}[X^{2}]}=\sqrt{\mathrm{Var}(X)}=\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (24)
Combining Eq. (23) and Eq. (24) gives the following guarantee under the linearized sensitivity:
$\epsilon_{\text{ada}}\approx\mathbb{E}|X|\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (25)
@@
Appendix B Additional ablation studies
B-A Computational Overhead of Prior Sensitivity under Larger Backbones
In addition to the pre-trained trackers used in this work (e.g., OSTrack-B256 with a ViT-B backbone), we further investigate the FLOPs and memory costs of the eigen-decomposition-based prior sensitivity estimation under larger backbones, such as ViT-L and Swin-L. Based on empirical measurements conducted on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, the specific computational overhead introduced by FIM estimation for ViT-L and Swin-L is detailed in Tab. XIV. It is important to note that the total cost of the offline FIM process is data-dependent, scaling with the size of the pre-training dataset and the number of iterations required for convergence given the model capacity. Consequently, rather than providing a variable total estimate, we report the deterministic per-group and per-iteration costs to accurately reflect the computational costs.
@@ B-A Computational Overhead of Prior Sensitivity under Larger Backbones
TABLE XIV: Computational cost of FIM eigen-decomposition on ViT-L and Swin-L ($\times256$ input resolution).
@@
B-B Training Efficiency of SRFT
@@
1. Accelerated Convergence: The FIM estimation is a one-shot process performed on the pre-training datasets. The resulting sensitivity map is then frozen and reused for all subsequent multi-modal tracking tasks. Furthermore, the pre-trained weight acts as a “smart initialization” compared to the random initialization of PEFT. By guiding the optimization along the most reliable directions from the start, SRFT significantly accelerates convergence. As shown in Tab. XV and Fig. 11, SRFT requires fewer optimal iterations and epochs to reach optimal performance compared to PEFT-based methods, effectively offsetting the preprocessing time and potentially achieving a lower total time-to-result for the full training lifecycle.
@@
3. Superior Performance: Finally, the computational cost translates directly into performance gains. By leveraging the global sensitivity information, SRFT consistently outperforms PEFT methods in accuracy (as detailed in the main paper tables), proving that the offline overhead yields a high return on performance.
TABLE XV: Comparison of preprocess time, trainable parameter and training time across different multi-modal tracking methods. All experiments are conducted based on the pre-trained OSTrack-B256. “Preprocessing” refers to prior sensitivity estimation. The training epochs (listed sequentially from left to right) correspond to the FE108, VisEvent, CoeSot, LasHeR, and DepthTrack datasets, respectively. Note that UnTrack employs a joint training strategy on the VisEvent, LasHeR, and DepthTrack datasets.
@@
Figure 12: Schematic illustration of different sensitivity fusion scheduling strategies. We compare the evolution of the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) using Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
B-C Different Sensitivity Fusion Scheduling Strategies
Our design is motivated by the stability–plasticity trade-off in cross-domain fine-tuning: effective transfer requires balancing stability (preserving critical pre-trained knowledge) and plasticity (adapting to the target domain). In the paper, we explicitly define a dynamic linear schedule to harmonize the two sensitivities: at the beginning, the prior sensitivity has weight $\kappa$, and as training proceeds the transfer sensitivity gradually increases until they reach the same weight $\kappa$ at the end. This is formalized by Eq. (14), where the combined sensitivity is a convex combination with linear dependence on training progress $\frac{t}{T}$.
As summarized in the Fig. 12 and Tab. XVI, we additionally compare exponential, cosine, piecewise (4-stage), and linear (ours) schedules on FE108 under sensitivity harmony coefficient $\kappa\in\{0.3,0.6,0.9\}$. To ensure a fair comparison of the trajectory effect induced by different schedules, we enforce the same initial and final weights for all schedules, so that the only difference lies in how the weights evolve over training. Overall, the linear schedule is consistently best or tied-best. When $\kappa=0.6$ (our default), both the weight evolution and the final performance are very close across different schedules, indicating the schedule choice is not overly sensitive in this moderate regime. However, when $\kappa$ becomes too small or too large, the alternative schedules tend to be more fragile. Mechanistically, exponential/cosine may over-emphasize one sensitivity at certain phases (e.g., too aggressive early transfer or overly conservative late adaptation), whereas piecewise scheduling introduces abrupt weight jumps that can destabilize optimization. In contrast, the linear schedule provides a monotonic and smooth transition with a constant rate of change, which empirically leads to more stable and robust behavior when $\kappa$ deviates from the moderate setting.
TABLE XVI: Comparison of different sensitivity fusion schedules on FE108 dataset. We compare the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) with Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
Schedule Prior Sensitivity Weight Transfer Sensitivity Weight $\kappa$ SR PR

### v3

@@ B.3 Derivation of Transfer Significance
Let $\mathcal{G}=\nabla_{\theta}\mathcal{L}(\theta\mid M)$ and recall:
@@
We model perturbations around a single update step as:
@@
Assume $L(\theta\mid M)$ is differentiable and locally $\beta$-smooth around $\theta$, i.e., $\|\nabla\mathcal{L}(\theta+\delta\mid M)-\nabla\mathcal{L}(\theta\mid M)\|_{2}\leq\beta\|\delta\|_{2}$. Then a first-order Taylor expansion yields:
@@
and hence:
$S(\theta,\delta\mid M)=-\mathcal{G}^{\top}\delta-r(\delta),$ (21)
Therefore, we obtain:
$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)=-\mathcal{G}^{\top}(\delta-\delta^{\prime})-(r(\delta)-r(\delta^{\prime})),$ (22)
When perturbations are sufficiently small (so the remainder term is dominated), the leading term is:
$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)\approx\mathcal{G}^{\top}(\delta^{\prime}-\delta)=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi),$ (23)
@@
Let $X:=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi)$. Since $\xi^{\prime}-\xi\sim\mathcal{N}(0,2\sigma^{2}I)$, we have $X\sim\mathcal{N}\!\big(0,\,2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}\big)$ and thus,
$\mathbb{E}|X|\leq\sqrt{\mathbb{E}[X^{2}]}=\sqrt{\mathrm{Var}(X)}=\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2},$ (24)
Combining Eq. (23) and Eq. (24) gives the following guarantee under the linearized significance:
$\epsilon_{\text{ada}}\approx\mathbb{E}|X|\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2},$ (25)
@@
Appendix C Additional Ablation Studies
C.1 Comparison with Sensitivity-aware Sparse Tuning (SPT).
In our method, parameter significance is used to regularize the update of all parameters, rather than to select a subset of “sensitive” parameters for sparse fine-tuning as in SPT (He et al., 2023). This distinction is particularly important for cross-modal adaptation. To examine this, we compare our approach with SPT under identical training settings, varying only the trainable parameter ratio $\tau$ for SPT (top-$\tau$ sensitive parameters).
As reported in Tab. 13, SPT exhibits a clear performance–sparsity trade-off: when $\tau$ decreases, tracking accuracy drops sharply (e.g., F-score $61.0\rightarrow 60.2\rightarrow 54.5$ for $\tau=50\%,20\%,10\%$), indicating that cross-modal adaptation requires distributed parameter updates rather than a highly sparse subset. Even with a relatively large budget ($\tau=50\%$), SPT remains essentially on par with full fine-tuning (F-score $61.0$ vs. $61.6$), suggesting that merely keeping the “most sensitive” parameters does not yield a meaningful adaptation gain beyond standard tuning. This observation is consistent with the nature of multi-modal tracking: the adaptation signal is typically heterogeneous across layers and operators, and restricting updates to a subset can easily break the coordinated adjustments needed for modality fusion and temporal consistency. In contrast, our method achieves a substantially higher F-score ($65.1$) while keeping $\tau=100\%$ trainable parameters, demonstrating that the key factor is not sparsifying the update but steering it. By reweighting parameter changes according to their significance, our regularization encourages the model to preserve parameters that are critical to general tracking behavior while allowing more flexible updates on parameters that are more transferable for cross-modal adaptation. As a result, our method improves both precision and recall (Pr $64.7$, Re $65.4$), indicating more accurate target localization and more complete target recovery under depth-induced appearance variations.
Table 13: Comparison of the tracking performance between SPT and our regularized fine-tuning method based on the DepthTrack dataset, using the pre-trained OSTrack-B256 as the base model. We have set up a series of trainable parameter ratios $\tau$ of SPT (top-$\tau$ sensitive parameters) to fully explore its adaption effect.
Exp. Pr Re F-score
w/o fine-tuning ($\tau=0\%$) 38.2 36.0 37.1
full fine-tuning ($\tau=100\%$) 61.7 61.5 61.6
SPT ($\tau=50\%$) 61.1 60.9 61.0
SPT ($\tau=20\%$) 60.7 59.7 60.2
SPT ($\tau=10\%$) 56.2 52.8 54.5
Ours ($\tau=100\%$) 64.7 65.4 65.1
@@ Appendix C Additional Ablation Studies
Table 14: Computational cost of FIM eigen-decomposition on OSTrack-B256 ($\times128$$\times256$ input resolution for template and search regions).
Metric Granularity Computational Cost
Per-Group (Operation-wise) Per-Iteration (Model-wise) Total Offline Process
FLOPs Theoretical Ops 104.8 GFLOPs 15.5 TFLOPs $\sim$ 39.7 EFLOPs
Latency Wall-clock Time $294.5$ ms $43.0$ s 47.8 h ($\sim$2 day)
[…]
Estimated over the full pre-trained tracking datasets. The preprocessing is performed only once.
C.2 Computational Overhead of Prior Significance Estimation
To theoretically justify the efficiency of prior significance estimation, we compare our approximation with the standard full-matrix estimation. Calculating the exact FIM on the global parameter space ($P\approx 86\text{M}$ for ViT-B) is computationally intractable, carrying a time complexity of $\mathcal{O}(P^{3})$ and a space requirement of $\mathcal{O}(P^{2})$ (theoretically requiring Petabytes of memory). In stark contrast, our SRFT adopts a group-wise approximation (operation-diagonal FIM) combined with Rayleigh quotient probing ($N_{\text{ops}}=148$ and $d=768$ for ViT-B). This strategy reduces the time complexity to $\mathcal{O}(N_{\text{ops}}\cdot d^{3})$—scaling linearly with the number of operations—and strictly bounds the peak memory to $\mathcal{O}(d^{2})$. Given that the operation width $d$ is magnitudes smaller than the total parameters $P$ ($d\ll P$), our method effectively reduces the computational cost from intractable to a practical level. We further profile this procedure on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, reporting theoretical FLOPs, latency, and peak GPU memory in Table 14 to quantify its efficiency. While this cost is justified by the resulting performance gains, we explicitly acknowledge that—unlike plug-and-play PEFT methods—prior significance estimation introduces additional, non-trivial computation. Importantly, it is a one-shot, offline pre-processing step performed only once before fine-tuning, and thus incurs negligible overhead during subsequent training iterations and inference. Moreover, it can accelerate convergence, effectively offsetting the preprocessing time and potentially reducing the overall time-to-result across the full training lifecycle.
C.3 Computational Overhead of Prior Significance under Larger Backbones
In addition to the pre-trained trackers used in this work (e.g., OSTrack-B256 with a ViT-B backbone), we further investigate the FLOPs and memory costs of the eigen-decomposition-based prior significance estimation under larger backbones, such as ViT-L and Swin-L. Based on empirical measurements conducted on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, the specific computational overhead introduced by FIM estimation for ViT-L and Swin-L is detailed in Table 15. It is important to note that the total cost of the offline FIM process is data-dependent, scaling with the size of the pre-training dataset and the number of iterations required for convergence given the model capacity. Consequently, rather than providing a variable total estimate, we report the deterministic per-group and per-iteration costs to accurately reflect the computational costs.
Table 15: Computational cost of FIM eigen-decomposition on ViT-L and Swin-L ($\times256$ input resolution).
@@ C.3 Computational Overhead of Prior Significance under Larger Backbones
C.4 Training Efficiency of Significance-Regularized Fine-Tuning
@@
1. Accelerated Convergence: The FIM estimation is a one-shot process performed on the pre-training datasets. The resulting significance map is then frozen and reused for all subsequent multi-modal tracking tasks. Furthermore, the pre-trained weight acts as a “smart initialization” compared to the random initialization of PEFT. By guiding the optimization along the most reliable directions from the start, SRFT significantly accelerates convergence. As shown in Tab. 16 and Fig. 12, SRFT requires fewer optimal iterations and epochs to reach optimal performance compared to PEFT-based methods, effectively offsetting the preprocessing time and potentially achieving a lower total time-to-result for the full training lifecycle.
@@
3. Superior Performance: Finally, the computational cost translates directly into performance gains. By leveraging the global significance information, SRFT consistently outperforms PEFT methods in accuracy (as detailed in the main paper tables), proving that the offline overhead yields a high return on performance.
Table 16: Comparison of preprocess time, trainable parameter and training time across different multi-modal tracking methods. All experiments are conducted based on the pre-trained OSTrack-B256. “Preprocessing” refers to prior significance estimation. The training epochs (listed sequentially from left to right) correspond to the FE108, VisEvent, CoeSot, LasHeR, and DepthTrack datasets, respectively. Note that UnTrack employs a joint training strategy on the VisEvent, LasHeR, and DepthTrack datasets.
@@
C.5 Different Significance Fusion Scheduling Strategies
Our design is motivated by the stability–plasticity trade-off in cross-domain fine-tuning: effective transfer requires balancing stability (preserving critical pre-trained knowledge) and plasticity (adapting to the target domain). In the paper, we explicitly define a dynamic linear schedule to harmonize the two significance: at the beginning, the prior significance has weight $\kappa$, and as training proceeds the transfer significance gradually increases until they reach the same weight $\kappa$ at the end. This is formalized by Eq. (14), where the combined significance is a convex combination with linear dependence on training progress $\frac{t}{T}$.
As summarized in the Fig. 13 and Tab. 17, we additionally compare exponential, cosine, piecewise (4-stage), and linear (ours) schedules on FE108 under significance harmony coefficient $\kappa\in\{0.3,0.6,0.9\}$. To ensure a fair comparison of the trajectory effect induced by different schedules, we enforce the same initial and final weights for all schedules, so that the only difference lies in how the weights evolve over training. Overall, the linear schedule is consistently best or tied-best. When $\kappa=0.6$ (our default), both the weight evolution and the final performance are very close across different schedules, indicating the schedule choice is not overly sensitive in this moderate regime. However, when $\kappa$ becomes too small or too large, the alternative schedules tend to be more fragile. Mechanistically, exponential/cosine may over-emphasize one significance at certain phases (e.g., too aggressive early transfer or overly conservative late adaptation), whereas piecewise scheduling introduces abrupt weight jumps that can destabilize optimization. In contrast, the linear schedule provides a monotonic and smooth transition with a constant rate of change, which empirically leads to more stable and robust behavior when $\kappa$ deviates from the moderate setting.
@@ C.5 Different Significance Fusion Scheduling Strategies
Figure 13: Schematic illustration of different significance fusion scheduling strategies. We compare the evolution of the prior significance fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) using Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial significance harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
Table 17: Comparison of different significance fusion schedules on FE108 dataset. We compare the prior significance fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) with Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial significance harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
Schedule Prior Significance Weight Transfer Significance Weight $\kappa$ SR PR